### v1

# Feedback-driven quantum reservoir computing for time-series analysis

Abstract
Quantum reservoir computing (QRC) is a highly promising computational paradigm that leverages quantum systems as a computational resource for nonlinear information processing. While its application to time-series analysis is eagerly anticipated, prevailing approaches suffer from the collapse of the quantum state upon measurement, resulting in the erasure of temporal input memories. Neither repeated initializations nor weak measurements offer a fundamental solution, as the former escalates time complexity while the latter restricts the information extraction from the Hilbert space. To address this issue, we propose a feedback-driven QRC framework. This methodology employs projective measurements on all qubits for unrestricted access to the quantum state, with the measurement outcomes subsequently fed back into the reservoir to restore the memory of prior inputs. We demonstrate that our QRC successfully acquires the fading-memory property through the feedback, a critical element in time-series processing. Notably, analysis of measurement trajectories reveal three distinct phases depending on the feedback strength, with the memory performance maximized at the edge of chaos. We also evaluate the predictive capabilities of our QRC, demonstrating its suitability for forecasting signals originating from quantum spin systems.
I INTRODUCTION
In the modern information society, quantum machine learning is gaining heightened interest for its potential in advanced data processing [1, 2, 3, 4, 5, 6, 7, 8]. The cornerstone of the advantage lies in the use of the exponentially large Hilbert space, which allows for the accommodation of high-dimensional features of datasets and the emulation of a substantial number of computational neurons [9, 10, 11, 12]. A quantum extension of the reservoir computing paradigm, known as quantum reservoir computing (QRC), also leverages this high dimensionality [13, 14]. Central to the philosophy of (general) reservoir computing is its distinctive architecture, where a fixed, randomly connected network of nodes–“reservoir”–processes input data by projecting it onto an internal feature space [15, 16, 17, 18, 19, 20, 21]. This strategy limits the training parameters only to the weights on the read-out part of the network, significantly reducing optimization costs and enabling high-speed, real-time temporal processing. Previous investigations have demonstrated that computational capabilities tend to be maximized when the reservoir part resides proximate to the boundary separating a stable and an unstable dynamical regimes, commonly referred to as the edge of chaos [22, 23, 24, 25, 26, 27]. Importantly, given the fixed nature of the internal parameters, the reservoir part can be realized either with an artificial recurrent neural network or with physical systems. In the QRC framework, a quantum system is harnessed as the physical embodiment of the reservoir part, termed a quantum reservoir. The intrinsic nonlinearity and high dimensionality of quantum systems perfectly align with the heuristic requirements for effective reservoir computing systems, which facilitate the extraction and transformation of intricate data features [13, 14, 19]. Thus far, theoretical proposals of the QRC have yielded promising results in addressing temporal and classification tasks [28, 29, 30, 31, 32, 33], and experimental demonstrations have been reported on several quantum reservoir settings, including nuclear magnetic resonance systems [34] and superconducting qubits [35]. Furthermore, since the information processing capabilities of the QRC are fundamentally linked to the nature of the quantum reservoirs [36, 37], it can be employed as a tool to investigate the quantum system itself, a concept recently formalized as quantum reservoir probing [38, 39].
Particularly focusing on temporal information processing, the fading-memory property emerges as a crucial aspect of the reservoir computing framework. This property dictates that when the reservoir projects input data onto the feature space, the temporal dependencies between different inputs are faithfully mapped within that internal space, and the influences of earlier inputs, being less significant than more recent ones, gradually diminish [15, 16]. Herein lies a significant challenge for the QRC: the inherent conflict between information extraction and memory preservation within the quantum reservoir. Conventional QRC architectures rely on quantum measurements to extract information from the quantum reservoir. However, these measurements inevitably induce measurement back-action, altering the quantum state of the reservoir. As a result, the memories of previous inputs encoded within the quantum reservoir are completely erased. This disturbance compromises the preservation of short-term memory and thus hinders the ability to process temporal information due to the absence of the crucial fading-memory property.
Figure 1: Architectures of QRC with an input sequence $\{s_{k}\}$ and an initial state $|\psi_{0}\rangle$. $\mathcal{L}$ symbolizes a generalized operation for a single input cycle parameterized by external inputs, the specific form of which depends on each realization. (a) Quadratic restarting protocol where the entire input signal is provided at every cycles to reinitialize the reservoir after the measurements. (b) Linear restarting protocol where the last $\tau$ inputs are repeated, assuming less effects from further previous inputs. (c) Online protocol that continuously monitors the time evolution using weak measurements or projective measurements only on limitied part of the system. (d) Feedback protocol where the measurement result at the $(k-1)$-th cycle determines the operation $\mathcal{L}$ at the $k$-th cycle.
To address this monitoring issue, a common strategy involves repeated initialization after each measurement. The entire process is executed from the beginning for every individual input, thereby effectively neglecting the effects of measurement back-actions [Fig. 1(a)] [35, 40, 41, 42]. However, this approach naturally necessitates $O(L^{2})$ input operations to address a time-series of length $L$, leading to a computationally expensive quadratic growth in execution time. A potential reduction in temporal complexity could be achieved under the premise of the fading-memory property. As the reservoir gradually loses information about distant past inputs, the simplification of the process by limiting input repetitions to recent cycles is expected to have minimal impact on memory performance [Fig. 1(b)]. By initiating the dynamics for the input cycle $k$ from the cycle $k-\tau+1$ rather than the cycle $1$, the total number of input repetitions reduces to $\tau L$, yielding a time complexity of $O(L)$ [43]. Nevertheless, despite this improvement, the requirement for repeating gate operations to input identical data still significantly compromises the real-time processing capabilities, a key advantage of the reservoir computing paradigm.
A promising approach seeks to balance short-term memory preservation with real-time functionality, by continuously monitoring quantum dynamics online while partially retaining the reservoir’s quantum state at each measurement [Fig. 1(c)]. By employing weak measurements or projective measurements on a limited subset of qubits, the QRC generates output online while mitigating the erasure of past input memory [44, 45, 46, 47, 48]. However, these approaches only permit limited information extraction, as partial measurements restrict comprehensive access to the quantum state of the reservoir. This constraint hinders the full utilization of the exponentially large Hilbert space in which the quantum reservoir resides, thus undermining a fundamental advantage of the QRC frameworks. Additionally, the necessity for supplementary ancillary systems in weak measurement configurations increases overall system complexity. Furthermore, considering the coherence time, there would be an upper bound on the length $L$ of the input data for which the system can maintain online. Consequently, the development of a simple and versatile online QRC protocol that fully leverages the quantum reservoir systems remains an urgent area of research.
In this paper, we propose a feedback-driven online QRC protocol that achieves complete access to the quantum state through projective measurements on all qubits [Fig. 1(d)]. While these measurements destroy short-term memory within the reservoir’s state, information of prior inputs is effectively revived by feedback connections that incorporate the measurement outcomes back into the quantum reservoir [49, 50, 51, 52]. By evaluating the short-term memory performance, we demonstrate that our QRC framework successfully acquire the fading-memory property through the feedback, showcasing its applicability to time-series analysis. Notably, the proposed QRC system is highly tunable, as the strength of input and feedback, which strongly relates to information embedding and memory preservation, can be controlled externally. Indeed, analysis of the measurement trajectories reveals three distinct phases depending on the feedback strength. The memory performance peaks at the phase boundary between the stable and the unstable phase, i.e., the edge of chaos. We further evaluate the predictive performance of our QRC for various time-series signals. Benchmark analysis against a conventional classical reservoir computing framework reveals that our QRC is highly suitable for forecasting signals derived from quantum spin systems. Our findings highlight the substantial potential of the QRC in the burgeoning quantum era as a potent instrument for time-series analysis.
The paper is organized as follows. In Sec. II, we introduce the framework of our feedback-driven QRC scheme. In Sec. III, we evaluate the short-term memory performance of the QRC. An in-depth analysis of the measurement trajectories, including the identification of the edge of chaos, is presented therein. In Sec. IV, we examine the prediction performance for the classical and quantum time-series signals. Section V is devoted to discussion and conclusion.
II METHODS
II.1 Architecture of the feedback-driven QRC
Figure 2(a) illustrates the architecture of our feedback-driven QRC system. We consider an $n$-qubit system initialized in the state $|\psi\rangle_{0}=|0\rangle^{\otimes n}$; in our numerical experiment, we specifically set $n=8$. Our QRC protocol consists of four steps: (i) input operation, (ii) feedback of previous measurement results, (iii) reservoir dynamics, and (iv) projective measurement on all qubits.
In step (i) of the $k$-th cycle, the input value $s_{k}$ is injected into the reservoir via the application of a two-qubit gate $\mathcal{R}_{1,2}(a_{\mathrm{in}}s_{k})$ on qubits $1$ and $2$, where $a_{\mathrm{in}}$ denote an input scaling weight. The operator $\mathcal{R}_{i,j}$ is defined as
$\mathcal{R}_{i,j}(\theta)=\mathrm{CX}_{ij}\mathrm{RZ}_{j}(\theta)\mathrm{CX}_{ij}\mathrm{RX}_{i}(\theta)\mathrm{RX}_{j}(\theta),$ (1)
which is commonly employed for embedding information in the QRC protocols [40, 41, 42]. Here, $\mathrm{CX}_{ij}$ denotes the CNOT gate acting on control qubit $i$ and target qubit $j$, while $\mathrm{RZ}(\theta)$ and $\mathrm{RX}(\theta)$ represent the rotation gate around the Pauli $Z$ and $X$ axis with angle $\theta$, respectively. The circuit diagram for $\mathcal{R}_{i,j}(\theta)$ is depicted in Fig. 2(b).
For clarity, we next describe the step (iv) of the ($k-1$)-th cycle. Here, Pauli $Z$ measurements are executed on all qubits, which yield the measurement result vector $\bm{z}_{k-1}=[\langle Z_{1}\rangle_{k-1},\langle Z_{2}\rangle_{k-1},\dots,\langle Z_{n}\rangle_{k-1}]^{\top}$ (where $Z_{i}$ denotes the Pauli $Z$ for the qubit $i$). Returning to step (ii) of the $k$-th cycle, the vector $\bm{z}_{k-1}$ is provided back into the reservoir through the feedback connection. The operator $\mathcal{R}_{i,j}(a_{\mathrm{fb}}z_{k-1}^{\alpha})$ is employed for feedback with weight $a_{\mathrm{fb}}$, which encodes all components of $\bm{z}_{k-1}$ according to the configuration shown in Fig. 2(a). Although the projective measurement strongly perturbs the quantum state, the feedback connections operate to systematically restore the memory of the provided input history.
Subsequent to the input and feedback processes, a fixed eight-qubit gate $U_{\mathrm{res}}$ is applied in step (iii) to generate entanglement among all qubits. This gate serves as the reservoir component of our QRC architecture, which simultaneously transforms the input and feedback information in a non-linear manner. We employ a Haar random unitary matrix for $U_{\mathrm{res}}$ in the following calculations. Finally, step (iv) concludes the $k$-th cycle with projective measurements on all qubits. The measurement result $\bm{z}_{k}$ is further cycled back to the reservoir via the feedback gate for processing the subsequent $(k+1)$-th input.
Figure 2: (a) Circuit diagram of the feedback-driven QRC architecture at the $k$-th cycle. The red and blue background color represent the gates for the input [step (i)] and feedback [step (ii)] parts, respectively. The input value $s_{k}$ is provided to qubits $1$ and $2$ via $\mathcal{R}_{1,2}(a_{\mathrm{in}}s_{k})$ with the input weight $a_{\mathrm{in}}$. The feedback of the measurement result $\bm{z}_{k-1}$ is given to qubits $3$ to $8$ via $\mathcal{R}_{i,j}(a_{\mathrm{fb}}z_{k-1}^{\alpha})$. Subsequently, the entanglement among all qubits is introduced by the reservoir gate $U_{\mathrm{res}}$ [step (iii)], and all qubits are then measured in the projective manner at the end of the circuit [step (iv)]. (b) Circuit representation of the two-qubit operator $\mathcal{R}$ defined in Eq. (1).
It is worth highlighting a key advantage of our architecture in terms of time complexity, which arises from the intrinsic independence of quantum circuits over different cycles. In our QRC, the quantum state at the $k$-th cycle depends exclusively on the current input value $s_{k}$ and the preceding measurement outcome $\bm{z}_{k-1}$, explicitly represented as $|\psi(s_{k},\bm{z}_{k-1})\rangle$. Even after the measurement, this state can be efficiently reconstructed by applying a limited set of gates, leveraging $\bm{z}_{k-1}$ stored in external memory; there is no need to re-execute the $(k-1)$-th and prior cycles when addressing the $k$-th input. Consequently, our protocol offers remarkable operational flexibility, enabling features such as starting from an intermediate cycle or running different cycles on different devices. In addition, owing to the frequent measurement and subsequent initialization, a coherence time sufficient for the execution of a single cycle is adequate, regardless of the length of the input data. These aspects stand in stark contrast to other implementations depicted in Figs. 1(a)-(c), which, due to their dependence on the input sequence as $|\psi(s_{k},s_{k-1},s_{k-2}\dots)\rangle$, necessitate significantly longer, uninterrupted operations.
II.2 Learning procedure
In the QRC protocol, the final output $y_{k}$ at the $k$-th cycle is obtained through a linear transformation of the measurement result $\bm{z}_{k}$. Here, a specific task for the QRC is formulated by setting a target output $\bar{y}_{k}$ based on the input sequence $\{s_{k}\}$. Our objective in the output part is to construct a regression model from the collection of measurement outcomes that accurately approximates the target output over various cycles. As a prerequisite, we define $\bm{z}_{0}$, required for the first cycle ($k=1$), as a random vector with each element sampled uniformly from the interval $[0,1]$. The influence of this initial condition is disregarded by discarding the initial $l_{\mathrm{w}}$ cycles.
The linear regression model is optimized for a training dataset containing $l_{\mathrm{tr}}$ inputs, whose performance is subsequently evaluated for unseen data in the testing phase for $l_{\mathrm{ts}}$ inputs. We define an $[l_{\mathrm{tr}}\times(n+1)]$-dimensional training matrix ${X}_{\mathrm{tr}}$ from the measurement results as
${X}_{\mathrm{tr}}=[{\bm{z}^{\prime}}_{l_{\mathrm{w}}+1},{\bm{z}^{\prime}}_{l_{\mathrm{w}}+2}\ldots{\bm{z}^{\prime}}_{l_{\mathrm{w}}+l_{\mathrm{tr}}}]^{\top},$ (2)
where $\bm{z}_{k}^{\prime}=\left[\bm{z}_{k}^{\top},1\right]^{\top}$ including an additional constant term. In the training phase, the weight vector $\bm{w}$ is optimized so that the output $\bm{y}_{\mathrm{tr}}={X}_{\mathrm{tr}}\bm{w}$ closely approximates the target value $\bar{\bm{y}}_{\mathrm{tr}}$. Utilizing the Moore-Penrose pseudo-inverse matrix, the optimal $\bm{w}$ is given by
$\bm{w}=(X^{\top}_{\mathrm{tr}}X_{\mathrm{tr}})^{-1}X^{\top}_{\mathrm{tr}}\bar{\bm{y}}_{\mathrm{tr}}.$ (3)
Following the training phase, we construct an $[l_{\mathrm{ts}}\times(n+1)]$-dimensional matrix ${X}_{\mathrm{ts}}$ for testing in a manner analogous to Eq. (2) using the measurement outcomes from the ($l_{\mathrm{w}}+l_{\mathrm{tr}}+1$)-th cycle to the ($l_{\mathrm{w}}+l_{\mathrm{tr}}+l_{\mathrm{ts}}$)-th cycle. The output is then calculated as $\bm{y}_{\mathrm{ts}}={X}_{\mathrm{ts}}\bm{w}$ using the trained weight. The similarity between the output $\bm{y}_{\mathrm{ts}}$ and the target output $\bar{\bm{y}}_{\mathrm{ts}}$ in the testing phase signifies the reservoir performance. As a quantitative metric, we employ the determination coefficient
$R^{2}=\frac{\mathrm{cov}(\bar{\bm{y}}_{\mathrm{ts}},\bm{y}_{\mathrm{ts}})}{\sigma^{2}(\bar{\bm{y}}_{\mathrm{ts}})\sigma^{2}(\bm{y}_{\mathrm{ts}})},$ (4)
where $\mathrm{cov}$ and $\sigma^{2}$ denote covariance and variance, respectively. $R^{2}$ approaches $1$ when the output $\bm{y}_{\mathrm{ts}}$ closely matches the target $\bar{\bm{y}}_{\mathrm{ts}}$; conversely, $R^{2}$ approaching 00 indicates a lack of correlation. We also employ the Normalized Mean Squared Error (NMSE):
$\mathrm{NMSE}=\frac{||\bar{\bm{y}}_{\mathrm{ts}}-\bm{y}_{\mathrm{ts}}||^{2}}{||\bar{\bm{y}}_{\mathrm{ts}}||^{2}},$ (5)
where |||||$\bullet$||| | ∙ | | denotes the Euclidean norm. The NMSE can assume any real value, and a low value of NMSE indicates a small deviation between the output and target. In our setup, we evaluate the reservoir performance with $l_{\mathrm{w}}=500$, $l_{\mathrm{tr}}=2000$, and $l_{\mathrm{ts}}=2000$.
III Fading-memory property with feedback connections
III.1 Short-term memory capacity
To evaluate the effectiveness of feedback connections for the preservation of input memories, we investigate the short-term memory task. This task involves processing an input sequence $\{s_{k}\}$ randomly sampled from a uniform distribution: $s_{k}\in[0,1]$. The goal is to reproduce the input value from $d$-cycle earlier, formulated as $\bar{y}_{k}=s_{k-d}$. The performance is measured using the determination coefficient $R^{2}_{d}$ for varying delays $d$, and the total capacity $C_{\Sigma}$ is defined by summing up these values
$C_{\Sigma}\equiv\sum_{d=0}^{d_{\mathrm{max}}}R^{2}_{d},$ (6)
where we set $d_{\mathrm{max}}=25$. Both $R^{2}_{d}$ and $C_{\Sigma}$ are averaged over $128$ samples with respect to the random matrix $U_{\mathrm{res}}$.
Figure 3: (a) Reservoir performance for the short-term memory task $R^{2}_{d}$ plotted as a function of the delay $d$ with $a_{\mathrm{fb}}=2.0$ (blue), $a_{\mathrm{fb}}=4.0$ (green), $a_{\mathrm{fb}}=6.0$ (purple), and $a_{\mathrm{fb}}=8.0$ (yellow). (b) The dependency of the total capacity $C_{\Sigma}$ on the feedback strength $a_{\mathrm{fb}}$. Both $R^{2}_{d}$ and $C_{\Sigma}$ are averaged over $128$ samples with respect to the $U_{\mathrm{res}}$.
Figure 3(a) exhibits the memory performance $R^{2}_{d}$ for varying feedback weights $a_{\mathrm{fb}}$, with the input weight fixed at $a_{\mathrm{in}}=0.001$. At a feedback strength of $a_{\mathrm{fb}}=2.0$, $R^{2}_{d}$ becomes close to $1$ for delays up to $d=4$, followed by a gradual decline for longer delays. Considering that the quantum state of the $k$-th cycle is determined only by $s_{k}$ and $\bm{z}_{k-1}$, this robust memory performance indicates that the input histories of $\{s_{k}\}$ are adequately accumulated within the measurement results $\bm{z}_{k-1}$. Essentially, the quantum reservoir effectively extracts key features of the provided dataset through the read-out measurements, and these features are succeeded over different cycles via the feedback connections. In this manner, our feedback-driven QRC overcomes the alternation of the quantum state caused by the measurement back-action, thereby acquiring the fading-memory property.
When the feedback is intensified to $a_{\mathrm{fb}}=4.0$, $R^{2}_{d}$ slightly diminishes for smaller delays ($d\leq 4$), as the influence of inputs becomes relatively subdued compared to the feedback effects. Conversely, for longer delays $d\geq 5$, $R^{2}_{d}$ shows an improvement characterized by a long-tailed structure as a result of the amplified feedback. The balance between the encoding and retention of the supplied information is thus tuned by $a_{\mathrm{fb}}$. Notably, excessively strong feedback destabilizes the system, evidenced by a decrease in $R^{2}_{d}$ at the feedback strength of $a_{\mathrm{fb}}=6.0$. Furthermore, at the even higher feedback strength ($a_{\mathrm{fb}}=8.0$), the rotation angle $a_{\mathrm{fb}}z^{i}_{k-1}$ in Eq. (1) exceeds $2\pi$, causing a malfunction of the feedback gates and resulting in $R^{2}_{d}\simeq 0$ for all $d$.
Figure 3(b) depicts the total capacity $C_{\Sigma}$ as a function of the feedback strength $a_{\mathrm{fb}}$, revealing three distinct phases: a stable phase for $0<a_{\mathrm{fb}}\lesssim 4.0$, an unstable phase for $4.0\lesssim a_{\mathrm{fb}}\lesssim 8.0$, and an over-rotation phase for $8.0\lesssim a_{\mathrm{fb}}$. In the stable phase, $C_{\Sigma}$ shows an increasing trend with $a_{\mathrm{fb}}$, which indicates that the stronger feedback more effectively revives older information. This phase represents a well-balanced state where the feedback mechanism enhances memory retention capabilities without excessively diminishing the influence of the new input. In contrast, the unstable phase is characterized by a decline in $C_{\Sigma}$ as feedback intensifies, due to the system being predominantly governed by feedback rather than input. Such feedback dominance compromises the reservoir’s functionality in processing and preserving information effectively. The over-rotation phase is marked by a complete collapse of $C_{\Sigma}$ to zero, triggered by an excessively large $a_{\mathrm{fb}}$ leading to an over-rotation upon the feedback. Consequently, optimal memory performance is achieved with feedback strength just below the threshold for the onset of unstable behavior [36]. This point, in close proximity to the boundary between the stable and unstable phases, corresponds to the edge of chaos in this system.
III.2 Measurement trajectories
Figure 4: Trajectories of the measurement results in (a), (c), (e) $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ plane and (b), (d), (f) $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$ plane for a typical instance of $U_{\mathrm{res}}$. Each point represents $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ at cycle $k$, and each line connect $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ and $\left[\langle Z_{i}\rangle_{k+1},\langle Z_{i+1}\rangle_{k+1}\right]$. We employ $a_{\mathrm{fb}}=3.0$ for (a) and (b), $a_{\mathrm{fb}}=6.0$ for (c) and (d), and $a_{\mathrm{fb}}=12.0$ for (e) and (f).
Since our QRC is driven by the feedback of the measurement outcomes and generates outputs as linear transformations thereof, an analysis of the measurement trajectories offers a deeper understanding into the physical origin underlying the three identified phases. Figure 4 displays the measurement results $\left[\langle Z_{1}\rangle_{k},\langle Z_{2}\rangle_{k}\right]$ and $\left[\langle Z_{3}\rangle_{k},\langle Z_{4}\rangle_{k}\right]$ at each cycle $k$ (similar trajectories are observed for $\langle Z_{5}\rangle_{k}$ to $\langle Z_{8}\rangle_{k}$). For $a_{\mathrm{fb}}=3.0$ in the stable phase, the dynamics of the quantum reservoir remain confined within a limited region of the feature space [Figs. 4(a) and 4(b)]. This behavior is attributable to the interplay between the feedback, acting to restore the pre-measurement state, and the input, introducing perturbations. The attractor-shaped trajectory within the feature space thus encapsulates the combined influences of both input and feedback. This encoding mechanism provides an understanding of the substantial preservation of short-term memories within the quantum reservoir in the stable phase.
However, this attractor-shaped dynamics collapses with further intensified feedback. Figures 4(c) and 4(d) illustrate typical dynamics in the unstable phase at $a_{\mathrm{fb}}=6.0$. Here, the dominant feedback causes the system to cycle through multiple stable states in the feature space, where the trajectory converges into a limit cycle. This undermines the influence of the input on the dynamics and, consequently, diminishes the system’s responsiveness to new inputs, leading to a decline in memory performance as $C_{\Sigma}\simeq 0$. Importantly, the formation of limit cycles depends on the specific realization of $U_{\mathrm{res}}$. For a particular $U_{\mathrm{res}}$ where the dynamics avoids entrapment in a limit cycle, attractor-like trajectories resembling those in Figs. 4(a) and 4(b) still emerge with a nonzero $C_{\Sigma}$ even at $a_{\mathrm{fb}}=6.0$. Nonetheless, as the feedback strength intensifies, the system exhibits an increased tendency to get caught in the limit cycle at a variety of $U_{\mathrm{res}}$ due to augmented instability. This results in a monotonic decrease in the capacity $C_{\Sigma}$ averaged over multiple instances of $U_{\mathrm{res}}$ in the unstable phase. In the over-rotation phase, the trajectory degrades into a random dynamics, with neither input nor feedback having a discernible influence [Figs. 4(e) and 4(f)]. These analysis underscore the significant influence of feedback connections on the dynamics of the quantum reservoir system, and thereby on the overall computational capabilities.
IV Prediction of classical and quantum time-series signals
IV.1 Dynamics of classical and quantum systems
Having demonstrated the preservation of the short-term memory facilitated by the feedback connections, we here evaluate the predictive capability of our QRC protocol for both classical and quantum time-series signals. This task is formulated by considering an input sequence $\{s_{k}\}$ normalized within the interval $[0,1]$, extracted from a specific dynamical system via time discretization. The aim of this task is to generate predictions $\tau_{f}$ cycles into the future, formally represented as $\bar{y}_{k}=s_{k+\tau_{f}}$. For performance evaluation of the QRC in predicting classical dynamics, we utilize the Mackey-Glass (MG) delay differential equation [16, 13, 53, 54]:
$\dot{x}(t)=\frac{\alpha x(t-t_{d})}{1+x(t-t_{d})^{\beta}}-\gamma x(t).$ (7)
With the parameters $\alpha=0.2$, $\beta=10$, and $\gamma=0.1$, the system manifests a chaotic attractor for $t_{d}>16.8$; our analysis centers on the MG dynamics with $t_{d}=17$. In formulating the input time-series signals, a single time-evolution step in the MG dynamics from $x(t)$ to $x(t+1)$ is represented by a single input cycle from $k$ to $k+1$ in the sequence $\{s_{k}\}$. In Fig. 5(a), the time-series is visually illustrated in the $s_{k}$-$s_{k+10}$ plane, along with an explicit representation as a function of time step $k$.
Additionally, to assess the efficacy in forecasting dynamics of quantum origin, we examine the spin dynamics of a 1D quantum Ising chain comprising five spins. The Hamiltonian is given by
$\mathcal{H}=-J\sum_{i}\sigma^{z}_{i}\sigma^{z}_{i+1}+h_{x}\sum_{i}\sigma^{x}_{i}+h_{z}\sum_{i}\sigma^{z}_{i}.$ (8)
Here, ${\sigma}_{i}^{x}$ and ${\sigma}_{i}^{z}$ denote the $x$ and $z$ Pauli matrices at the site $i$, while $h_{x}$ and $h_{z}$ signify the transverse and longitudinal magnetic fields, respectively. $J>0$ corresponds to the strength of the ferromagnetic nearest-neighbor Ising interaction, normalized to $J=1$ for our energy scale. Our focus lies in predicting the expectation value of the central spin’s $z$-component, $\langle\sigma^{z}_{3}(t)\rangle$, where the quantum system is initialized in the all up state and evolving according to the Heisenberg equation under the Hamiltonian in Eq. (8). In the absence of a longitudinal field ($h_{z}=0$), the spin system can be mapped to a free fermion system via the Jordan-Wigner transformation, leading to integrable dynamics [55]. Conversely, at $(h_{x},h_{z})=(1.05,-0.5)$, the system exhibits nonintegrable dynamics characterized by the chaotic spectral statistics [56]. The time-series $\{s_{k}\}$ is constructed by mapping the dynamics of the quantum spin from $t$ to $t+0.05$ to a single step from $k$ to $k+1$; visual representations of these inputs are provided in Fig. 5(a).
IV.2 Predictive performance of the QRC
Figure 5: (a) Visualization of the time-series $\{s_{k}\}$ derived from classical MG dynamics, integrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$, and nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.05,-0.5)$. Each $\{s_{k}\}$ is depicted in the $s_{k}$-$s_{k+10}$ plane (top) and as a function of the time step $k$ (bottom). (b)-(e) Performance of the QRC and the ESN for the predictions of (b) cosine wave, (c) classical MG time-series, (d) integrable quantum spin dynamics, and (e) nonintegrable quantum spin dynamics. NMSE is averaged for $128$ random realizations for both the QRC and ESN. The QRC with $8$ qubits is constructed at $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{back}}=2.5$, and the ESN is equipped with $8$, $50$, and $100$ nodes. The performance of the ESN represents the lowest NMSE achieved through the optimization of network parameters.
Figure 6: Measurement trajectories for a typical $U_{\mathrm{res}}$ in $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$, $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$, $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$, and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes. Each point represents $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ at cycle $k$, and each line connect $\left[\langle Z_{i}\rangle_{k},\langle Z_{i+1}\rangle_{k}\right]$ and $\left[\langle Z_{i}\rangle_{k+1},\langle Z_{i+1}\rangle_{k+1}\right]$. The input and feedback strength are set to $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{back}}=2.5$, respectively. The input sequence $\{s_{k}\}$ is derived from (a) cosine wave, (b) classical MG time-series, (c) integrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$, and (d) nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.05,-0.5)$.
In Figs. 5(b)-5(e), we present the predictive performance of our QRC across four different time-series: a cosine wave $\{s_{k}\}=\{\cos(\omega k)\}$ with $\omega=\pi/25$, the classical MG time-series with $t_{d}=17$, and both the integrable and nonintegrable quantum spin dynamics at $(h_{x},h_{z})=(1.0,0.0)$ and $(h_{x},h_{z})=(1.05,-0.5)$, respectively. We employ the input and feedback weights of $a_{\mathrm{in}}=0.001$ and $a_{\mathrm{fb}}=2.5$, ensuring that the system remains in the stable phase when subjected to random inputs (Figs. 3 and 4). For comparison, we present the performance of an echo state network (ESN), a typical classical reservoir computing framework, while varying the number of nodes and optimizing network parameters for each $\tau_{f}$; see Appendix A for further details.
In predicting the cosine wave, the QRC exhibits remarkable performance, surpassing or comparable to the $100$-node ESN; $\mathrm{NMSE}\simeq 10^{-9}$ is achieved across all $\tau_{f}$ [Fig. 5(b)]. Moreover, it demonstrates pronounced predictive accuracy in forecasting both integrable and nonintegrable quantum spin dynamics, outperforming the $100$-node ESN up to $\tau_{f}\simeq 15$ [Figs. 5(d) and 5(e)]. (For $\tau_{f}>15$, both the QRC and the ESN perform poorly with $\mathrm{NMSE}>10^{-3}$, making comparisons in this regime less meaningful.) In contrast, for predictions of the classical chaotic MG dynamics, the effectiveness of our QRC diminishes, falling below the performance of both the $50$-node and $100$-node ESN for $\tau_{f}>3$, as illustrated in Fig. 5(c). These results highlight the strengths and limitations of our QRC in predicting diverse categories of time-series data: it shows a distinct advantage in predicting time-series originating from quantum spin dynamics, yet is less adept at predicting classical MG dynamics.
Notably, our eight-qubit QRC consistently outperforms the ESN with an equivalent number of nodes and, for specific targets, even surpasses the $100$-node ESN. It is worth highlighting that our system necessitates the optimization of only $8+1$ parameters at the output part (including the bias term), which is a significant reduction compared to the $100+1$ optimization parameters in the $100$-node ESN. The remarkable performance of the QRC, despite having fewer computational nodes, can be attributed to the high-dimensionality of the quantum reservoir’s Hilbert space, widely recognized as a highly advantageous characteristic for effective reservoir computing [13, 19]. The reduction in computational nodes also contribute to the superior performance of the QRC at $\tau_{f}=0$; in such scenarios, where the objective is simply to reproduce the most recent input, excessive complexity can be detrimental.
Figure 6 illustrates the measurement trajectories under the influence of each time-series signal $\{s_{k}\}$; for a typical realization of $U_{\mathrm{res}}$, $\left[\langle Z_{1}\rangle_{k},\langle Z_{2}\rangle_{k}\right]$, $\left[\langle Z_{3}\rangle_{k},\langle Z_{4}\rangle_{k}\right]$, $\left[\langle Z_{5}\rangle_{k},\langle Z_{6}\rangle_{k}\right]$, and $\left[\langle Z_{7}\rangle_{k},\langle Z_{8}\rangle_{k}\right]$ are depicted. When subjected to the cosine wave input, the dynamics of the quantum reservoir system exhibits periodic behavior, manifested as elliptical orbital trajectories in the feature space [Fig. 6(a)]. These trajectories possess a periodicity of $50$ cycles, which is consistent with the period of the original cosine wave input. This observation indicates that the internal dynamics of the quantum reservoir adequately reflects the provided input dynamics.
In Fig. 6(b), we present the measurement outcomes influenced by the MG dynamics input. These trajectories, presumably reflecting the chaotic attractor observed in Fig. 5(a), form a complicated manifold that cannot be easily reduced to a simple two-dimensional representation. Conversely, under the input signals derived from integrable [Fig. 6(c)] and nonintegrable [Fig. 6(d)] quantum spin dynamics, the trajectories collapse into single lines in specific two-dimensional planes: the $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$, $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$, and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes in Fig. 6(c), and the $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ and $\langle Z_{3}\rangle$-$\langle Z_{4}\rangle$ planes in Fig. 6(d). Even when trajectories display broader dispersions, they predominantly manifest as superpositions of elliptical curves, each with varying radii and central locations [the $\langle Z_{1}\rangle$-$\langle Z_{2}\rangle$ plane in Fig. 6(c), and the $\langle Z_{5}\rangle$-$\langle Z_{6}\rangle$ and $\langle Z_{7}\rangle$-$\langle Z_{8}\rangle$ planes in Fig. 6(d)]. This rotation-type reservoir dynamics is attributed to the characteristics of the provided signals, which originate from quantum spin dynamics that fundamentally involves spin rotations.
Although the raw time-series signals presented in Fig. 5(a) alone do not reveal their suitability for the QRC-based prediction, trajectories associated with successful predictions (i.e., cosine and quantum spin dynamics) lead to the inference that input signals inducing rotation-like internal reservoir dynamics are suitable for prediction by the QRC. Interestingly, even among such signals, our QRC exhibits varying predictive performance for different time-series. For instance, its prediction performance for the nonintegrable quantum spin dynamics is somewhat diminished compared to that for the integrable dynamics, as depicted in Figs. 5(d) and 5(e). This discrepancy might be attributable to the increased complexity of the nonintegrable dynamics, corroborated by the observation that measurement trajectories in the integrable case often converge towards simpler patterns compared to the nonintegrable case. Further classification of time-series signals based on their predictability within the QRC framework will be a subject of future research.
V Discussion and conclusion
In summary, we have developed the feedback-driven QRC framework, which is characterized by the incorporation of feedback connections to the measurement outcomes. Similar to existing QRC models, our approach utilizes quantum measurements for information extraction from the quantum reservoir, wherein the quantum state is altered at each cycle. Nevertheless, we have demonstrated that memories of previous inputs are effectively reincorporated into the reservoir through feedback processes, which endows our QRC system with the fading-memory property. Notably, this framework offers distinct advantages over existing QRC models, particularly in the unrestricted access to the quantum state via projective measurements on all qubits, not limited to weak or partial measurements, without increasing time complexity. Moreover, we have elucidated that our QRC system manifests three distinct phases based on the feedback strength, as evidenced by the measurement trajectories. The memory performance is maximized when the feedback positions the quantum reservoir system in proximity to the boundary between stable and unstable phases, corresponding to the edge of chaos. In time-series predictions, we have shown that the QRC demonstrates proficiency in forecasting dynamics that induces rotations within the quantum reservoir; it outperforms the $100$-node ESN for specific prediction targets despite its limited number of optimization parameters.
Remarkably, the high tunability of our QRC would facilitate task-adaptive performance optimization [57]. Conventional physical reservoir systems possess few accessible hyperparameters, rendering their reservoir properties largely fixed and making task-dependent optimization a general challenge for reservoir computing frameworks. Our protocol will overcome this challenge by flexibly controlling the internal dynamics of the quantum reservoir through varying the feedback strength. Combined with its online feature, we anticipate further advancements enabling real-time, on-demand performance tuning through adaptive feedback adjustments. Notably, feedback mechanisms are prevalent in the field of quantum machine learning, as exemplified by variational quantum algorithms (VQAs) [2, 6, 7]. In VQAs, parameters for quantum variational circuits are iteratively trained based on feedback of a cost function calculated from measurement outcomes. Although the proposed QRC procedure does not involve variational optimizations, it shares common features with VQAs, particularly in the utilization of measurements and their subsequent feedback. Consequently, our QRC will be further enhanced by leveraging insights from VQAs, especially in feedback configurations.
We highlight the versatility of our feedback-driven QRC scheme, applicable to a broad spectrum of quantum reservoir platforms, encompassing both fermionic and bosonic systems [58, 59, 60, 61, 62]. The intrinsic physical properties of these platforms will lead to diverse information processing capabilities, particularly in the types of time-series that can be effectively predicted. This emphasizes the importance of investigating a wide array of potential quantum reservoir platforms. Notably, such extensive exploration will also provide insights into quantum many-body phenomena from unconventional perspectives, leveraging the effectiveness of the QRC to probe the quantum reservoir system itself [38, 39]. We anticipate that these investigations will catalyze advancements in both quantum information and quantum materials, thereby propelling the field of quantum information science forward.
Acknowledgements.
K.K. thanks Yukitoshi Motome for insightful discussions. The authors acknowledge the support from METI and IPA through the MITOU Target program. K.K. is supported by the Program for Leading Graduate Schools (MERIT-WINGS) and JSPS KAKENHI Grant Number JP24KJ0872. K.F. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0120319794, JST COI- NEXT Grant No. JPMJPF2014. N.Y. is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grants No. JPMXS0118067285 and No. JPMXS0120319794.
Appendix A Echo state network (ESN)
ESN is a well-established framework in classical reservoir computing [15, 16, 18]. For an ESN characterized by its number of nodes $N_{\mathrm{node}}$, the internal state at cycle $k$ is represented as an $N_{\mathrm{node}}$-dimensional vector $\bm{z}^{\mathrm{ESN}}_{k}$. This state is updated iteratively by
$\bm{z}^{\mathrm{ESN}}_{k+1}=(1-\mathrm{lr})\bm{z}^{\mathrm{ESN}}_{k}+\mathrm{lr}f(W_{\mathrm{in}}s_{k+1}+W\bm{z}^{\mathrm{ESN}}_{k}),$ (9)
where $\mathrm{lr}$ signifies the learning rate, and $f$ represents the element-wise activation function, for which we adopt $f(x)=\tanh(x)$. $W_{\mathrm{in}}$ represents the $N_{\mathrm{node}}$-dimensional input weights, each assigned values of either $1$ or $-1$ with an equal probability. $W$ denotes the $(N_{\mathrm{node}}\times N_{\mathrm{node}})$-dimensional recurrent weights matrix, whose elements are sampled from a Gaussian distribution and scaled to achieve a spectral radius of $\lambda$. These internal weights remain fixed throughout the process. In the learning phase, the training matrix $X$ is constructed using the state vector ${\bm{z}^{\mathrm{ESN}}_{k}}^{\prime}=[{\bm{z}^{\mathrm{ESN}}_{k}}^{\top},1]^{\top}$ in the same manner as Eq. (2). The output weight $\bm{w}$ is trained via Ridge regression with a regularization term $\beta=10^{-5}$ to prevent instability, expressed by
$\bm{w}=(X^{\top}X+\beta I)^{-1}X^{\top}\bar{\bm{y}}_{\mathrm{tr}}.$ (10)
The final output at the $k$-th cycle is computed as $y_{k}={\bm{z}^{\mathrm{ESN}}_{k}}^{\prime\top}\bm{w}$, and the performance is evaluated using Eq. (5). In Figs. 5(b)-5(e), we present the optimal performance of the ESN by varying the parameters within the ranges $\lambda\in[0.5,0.75,0.95,1.25,1.5]$ and $\mathrm{lr}\in[0.4,0.6,0.8,1.0]$.